OpenStudy (anonymous):

The Howard family and the Anderson family each used their sprinklers last summer. The water output rate for the Howard family's sprinkler was 30L per hour. The water output rate for the Anderson family's sprinkler was 15L per hour. The families used their sprinklers for a combined total of 35 hours, resulting in a total water output of 750L How long was each sprinkler used?

OpenStudy (anonymous):

You can make 2 equations, with 2 variables. 30 * x + 15 * y = 750 x + y = 35 Now you just solve x and y.

OpenStudy (anonymous):

Do you understand where the equations come from? and do you know how to solve it?

OpenStudy (anonymous):

yes but idk how to do the second part of the problem x+y=35

OpenStudy (anonymous):

You can solve 2 equations with 2 variables in many different ways, here is a simple way: Firs you isolate x in 1 of the equations: x + y = 35 <=> x = 35 - y, Now you insert 35 - y on x in the other equation, so you get: 30 * (35 - y) + 15 * y = 750, now you isolate y and you get y = 20. And now you can insert that in one of the equations: x + 20 = 35 <=> x = 15 So the answer is y = 20 and x = 15, Does that make sense?

OpenStudy (anonymous):

yes it does and is x the howards or is y the howards

OpenStudy (anonymous):

That method is called the system of linear equations. You can read some more about it here: https://en.wikipedia.org/wiki/System_of_linear_equations

OpenStudy (anonymous):

@Tommynaiter alright thanks but is x howards or is y howard

OpenStudy (anonymous):

Howard is x, because i made the equation 30 * x + 15 * y = 750, and howard used 30L/hour, and the other used 15 L/hour, and in total they used 750L

OpenStudy (anonymous):

And the second equation x + y = 35, is made by the fact that they used 35 hours in total

OpenStudy (anonymous):

okay thx